# Nonlocal Josephson diode effect in minimal Kitaev chains

### v1

@@ Abstract
The research on nonlocal JE has not only attracted theoretical efforts [10, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] but its realization has already been reported in many experiments [36, 37, 38, 39, 40, 41, 42]. Most of the studies involved conventional spin-singlet $s$-wave superconductors, leaving unconventional superconductors seldom explored in spite of their intrinsic nonlocal order parameters. One of the simplest systems that hosts controllable unconventional superconductivity with a $p$-wave order parameter can be engineered by connecting two quantum dots (QDs) via a superconductor [43, 44]. This setup represents a minimal Kitaev chain, where electron cotunneling (ECT) and crossed Andreev reflections (CARs) mediate superconducting and normal processes between QDs [43, 45]. Minimal Kitaev chains have lately triggered multiple efforts [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] and their implementation has already been shown in superconductor-semiconductor hybrids [59, 60, 61, 62]. Despite the significant attention, the nonlocal JE and its impact on minimal Kitaev chains have not yet been studied.
@@
JJs formed by three minimal Kitaev chains.—We consider a Josephson system where three minimal Kitaev chains described by
@@
Here $\alpha={\rm L,M,R}$ labels the left (L), middle (M), and right (R) minimal Kitaev chains formed by two sites $j=1,2$, and $c_{\alpha_{j}}$ ($c_{\alpha_{j}}^{\dagger}$) destroys (creates) an electronic state at site $j=1,2$ in the chain $\alpha$. The first and second terms in $H_{\alpha}(\phi_{\alpha})$ describe onsite energies $\varepsilon_{\alpha_{j}}$ controlled by gates [brown bars in Fig. 1]. The third term represents hopping between sites on the same chain, a characteristic of ECT processes. The fourth term is a spinless superconducting potential, with amplitude $\Delta_{\alpha}$ and phase $\phi_{\alpha}$, describing CAR processes. We model the coupling between the three chains by the second and third terms in Eq. (2) with tunneling amplitudes $\tau_{\rm L(R)}$ across the left (right) junctions. Using a gauge symmetry, we also set the phase of the middle chain as $\phi_{\rm M}=0$, leading to the junction depending on two superconducting phases, $H(\phi_{\rm L},\phi_{\rm R})$. The phase differences $\phi_{\rm L}$ and $\phi_{\rm R}$ between chains can be externally controlled by magnetic fluxes [Fig. 1], as in Refs. [40, 63, 42]. Despite the simplicity, two-site Kitaev chains have already been experimentally realized in QDs coupled by a superconductor-semiconductor system [59, 60, 61, 62, 64], where the control of ECT and CAR was demonstrated [65, 66]. As we show below, the system exhibits a unique phase dependence, enabling a nonlocal JE with nonreciprocal origin.
@@
Breaking and preservation of symmetries.—We begin by analyzing symmetries of the Josephson system in Eq. (2). First of all, the quasi-particle energy derived by Eq. (2) forms a pair (E,E)(-E,E) because of particle-hole symmetry (PHS) of the Bogoliubov-de Gennes Hamiltonian. While the Josephson system respects time-reversal symmetry (TRS), $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},-\phi_{\rm R})$, where the time-reversal anti-unitary operator $\cal T$} is given by ${\cal T}c_{\alpha_{i}}{\cal T}^{-1}=c_{\alpha_{i}}$, it does not have either $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},\phi_{\rm R})$ or $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(\phi_{\rm L},-\phi_{\rm R})$, implying local TRS is broken either across the left or right junctions, respectively. For $\tau_{\rm L(R)}\neq 0$, the left (right) unction also explicitly breaks the local spatial inversion symmetry, even when the entire system may be spatially inversion-symmetric. Thus, each junction satisfies the standard criterion for the Josephson diode effect [67, 68, 69]. However, our system Eq. (2) has an additional symmetry constraint for the Josephson diode effect, which originates from a special symmetry at the “sweet spot” ($\varepsilon_{\alpha}=0$, $t_{\alpha}=\Delta_{\alpha}$), where the chains support quadratically protected Majorana states [43]. In fact, for $\varepsilon_{\alpha}=0$ and $t_{\alpha}=\Delta_{\alpha}$, each minimal chain has a local charge-conjugation symmetry, ${\cal C}_{\alpha_{1}}H_{\alpha}(\phi_{\alpha}){\cal C}_{\alpha_{1}}^{-1}=H_{\alpha}(-\phi_{\alpha})$, where the non-trivial action of the unitary operator ${\cal C}_{\alpha_{1}}$ is defined by ${\cal C}_{\alpha_{1}}c_{\alpha_{1}}{\cal C}_{\alpha_{1}}^{-1}=c^{\dagger}_{\alpha_{1}}$, ${\cal C}_{\alpha_{1}}c_{\alpha_{2}}{\cal C}_{\alpha_{1}}^{-1}=-e^{i\phi_{\alpha}}c_{\alpha_{2}}$, ${\cal C}_{\alpha_{1}}c_{\beta_{i}}{\cal C}_{\alpha_{1}}^{-1}=c_{\beta_{i}}$ $(\beta\neq\alpha)$. Note that ${\cal C}_{\alpha_{1}}$ switches only an electron at site 1 on chain $\alpha$ with its antiparticle. We also have a similar local charge-conjugation symmetry ${\cal C}_{\alpha_{2}}$, switching only an electron at site 2 with its antiparticle. We can partially keep the symmetries even for the junctions in Eq. (2). For instance, by performing ${\cal C}_{{\rm L}_{2}}$, ${\cal C}_{{\rm M}_{1}}$, and an additional gauge rotation, we have ${\cal C}_{\rm LM}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm LM}=H(-\phi_{\rm L},\phi_{\rm R})$, where the unitary operator ${\cal C}_{\rm LM}$ acts as ${\cal C}_{\rm LM}c_{{\rm L}_{1}}{\cal C}^{-1}_{\rm LM}=e^{i\phi_{\rm L}}c_{{\rm L}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm L}_{2}}{\cal C}^{-1}_{\rm LM}=c^{\dagger}_{{\rm L}_{2}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{1}}{\cal C}_{\rm LM}^{-1}=-c^{\dagger}_{{\rm M}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{2}}{\cal C}_{\rm LM}^{-1}=c_{{\rm M}_{2}}$, and ${\cal C}_{\rm LM}c_{{\rm R}_{i}}{\cal C}_{\rm LM}^{-1}=c_{{\rm R}_{i}}$. We also have the relation ${\cal C}_{\rm MR}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm MR}=H(\phi_{\rm L},-\phi_{\rm R})$ with an unitary operator ${\cal C}_{\rm MR}$ constructed similarly. These relations provide the same constraint on the quasi-particle spectrum as local inversion, $E_{n}(\phi_{\rm L},\phi_{\rm R})=E_{n}(\pm\phi_{\rm L},\pm\phi_{\rm R})$, and thus, we need to break these local charge-conjugation symmetries, in addition to local TRS, to obtain the Josephson diode effect. For this purpose, we consider $\Delta_{\rm M}\neq t_{\rm M}$ below, which breaks both ${\cal C}_{\rm LM}$ and ${\cal C}_{\rm MR}$.
@@
Andreev bound states and Andreev molecules.—Having identified the role of the superconducting phase differences $\phi_{\rm L,R}$ on the system symmetries, here we explore how they determine the behavior of the Andreev energy spectrum. This is shown in Fig. 2(a-d) as a function of $\phi_{\rm L}$ when ECT and CAR in the middle chain are the same ($t_{\rm M}=\Delta_{\rm M}$) for distinct values of $\phi_{\rm R}$ at $\Delta_{\rm L,R}=t_{\rm L,R}$, while in Fig. 2(e-h) for ECT weaker than CAR. Since $\Delta_{\rm L,R}=t_{\rm L,R}$, the left (right) chains host zero-energy dispersionless Majorana flat bands located in their first (second) QD, which remain robust against $\phi_{\rm L,R}$ [Fig. 2]. Depending on the ratio between CAR and ECT in the middle chain, the Andreev spectrum exhibits a strong dependence on $\phi_{\rm L,R}$. For $t_{\rm M}=\Delta_{\rm M}$ at $\phi_{\rm R}=0$ [Fig. 2(a)], a pair of dispersionless Andreev bound states (ABSs) appear at energies close to $\pm\Delta$ and intersect a pair of dispersing ABSs near $\phi_{\rm L}=0,2\pi$. See green vertical lines. These dispersionless ABSs originate from the right junction between the middle and right chains, and, as $\phi_{\rm R}$ takes finite values, they hybridize with the dispersing ABSs formed in the left junction between the left and middle chains [Figs. 2(a-d)]. The dispersing hybridized ABSs, which can be seen as Andreev molecules studied before in conventional superconductors [10, 70, 71, 26, 72, 41, 73, 63, 33], reach zero energy at $\phi_{\rm L}=\pi$ (gray vertical line) and become flat bands with $\phi_{\rm L}$ at $\phi_{\rm R}=\pi$ [Fig. 2(c)], always maintaining a symmetric profile about $\phi_{\rm L}=\pi$. When $t_{\rm M}\neq\Delta_{\rm M}$, a nonzero $\phi_{\rm R}$ modifies the phase dependence of the spectrum and, notably, it also makes the spectrum asymmetric about $\phi_{\rm L}=\pi$ (vertical dashed red line), provided $\phi_{\rm R}\neq 2\pi n$ with $n\in\mathbb{Z}$; see Fig. 2(e-h) and Fig. 5 in S1 of the End Matter. This asymmetry moves the zero-energy crossing of the dispersing hybridized ABSs to $\phi_{\rm L}\neq\pi$, which occurs below (above) $\phi_{\rm L}=\pi$ when $\phi_{\rm R}<\pi$ ($\phi_{\rm R}>\pi$) for $t_{\rm M}<\Delta_{\rm M}$ [Fig. 2(e-h)]. The asymmetric spectrum turns out to be gapless over $\phi_{\rm L,R}\in(0,2\pi)$ and remains for $t_{\rm M}>\Delta_{\rm M}$ but with the zero-energy crossing at $\phi_{\rm L}<\pi$ ($\phi_{\rm L}>\pi$) when $\phi_{\rm R}>\pi$ ($\phi_{\rm R}<\pi$) in the latter case [Fig. 5]. A consequence of the asymmetric spectrum under $\phi_{\rm L,R}\neq 0$ and $\Delta_{\rm M}\neq t_{\rm M}$ is that the free energy develops minima away from $\phi_{\rm L}=2\pi n$, unlike conventional Josephson junctions [74, 75]; (See S1 in End Matter). Thus, the imbalance between CAR and ECT in the middle Kitaev chain induces an asymmetric Andreev spectrum at nonzero superconducting phases across the left/right junctions, in line with breaking local TR and local charge-conjugation symmetries.
@@
Local and nonlocal Josephson effect.—A direct consequence of $\phi_{\rm L,R}\neq 0$ and the imbalance between CAR and ECT in the middle chain is the phase-biased Josephson effect, which we explore here. For this purpose, we numerically obtain the phase-biased Josephson current flowing across the left junction as [76, 77, 24] $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm L}$  [^1] Similarly, we can obtain the supercurrent flowing across the right junction as $I_{\rm R}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm R}$. We have verified that $I_{\rm L,R}(\phi_{\rm L},\phi_{\rm R})$ can also be obtained within a Green’s function approach [102, 77], but obtaining them from the spectrum offers a direct way to identify the contributing states [103, 104, 105, 106, 107, 108, 109].. To visualize the supercurrent, in Figs. 3(a,c,e) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ and $\phi_{\rm R}$ for distinct cases when CAR in the middle chain is equal (a), weaker (c), and stronger (e) than ECT in the same chain. In Figs. 3(b,d,f) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ for distinct $\phi_{\rm R}$, while Figs. 3(d,f) also show $I_{\rm L}$ as a function of $\phi_{\rm R}$ for distinct $\phi_{\rm L}$. The first observation is that $I_{\rm L}$ in Figs. 3 strongly depends on $\phi_{\rm L,R}$ and the ratio between ECT ($t_{\rm M}$) and CAR ($\Delta_{\rm M}$) in the middle region, a behavior that originates from the ABSs in Figs. 2. When $\Delta_{\rm M}=t_{\rm M}$, the supercurrent $I_{\rm L}$ is $2\pi$-periodic and symmetric under $\phi_{\rm R}\neq 0$ [Figs. 3(a,b)], having $I_{\rm L}(\phi_{\rm L}=n\pi)=0$ for any $\phi_{\rm R}$ and equal maxima/minima that reduce for $\phi_{\rm R}\neq 0,2\pi$; the free energy in Fig. 6(a) has a minimum at $\phi_{\rm L}$ as for 0-junctions [74, 79]. Also, $I_{\rm L}$ exhibits a sawtooth profile at $\phi_{\rm L}=\pi$ for $\phi_{\rm R}\neq\pi$ due to the zero-energy crossing of the dispersing ABSs [Figs. 2(a-d)], while it smoothens out at $\phi_{\rm R}=\pi$ because the dispersing ABSs become flat [Figs. 2(c)].
@@
At $\Delta_{\rm M}\neq t_{\rm M}$, the supercurrent across the left junction $I_{\rm L}$ develops an intriguing anomalous behavior that varies with $\phi_{\rm L,R}\neq 0$. See Figs. 3(c-f). The most notorious feature is that $\phi_{\rm R}\neq n\pi$ gives rise to $I_{\rm L}(\phi_{\rm L}=n\pi)\neq 0$ with $n\in\mathbb{Z}$ and it is not symmetric about $\phi_{\rm L}=\pi$ anymore, which are consequences of the asymmetric Andreev spectrum shown in Figs. 2(e-h). This implies that, even at $\phi_{\rm L}=0$, a supercurrent across the left junction $I_{\rm L}$ emerges and is controlled by the superconducting phase difference across the right junction $\phi_{\rm R}$, signalling the so-called nonlocal JE which was initially predicted in systems with conventional superconductors [10]. The sawtooth profile of $I_{\rm L}$ also occurs away from $\phi_{\rm L}=\pi$ [Figs. 3(d,f)] due to the shifted zero-energy crossing in the ABSs when $t_{\rm M}\neq\Delta_{\rm M}$. See Figs. 2(e-h) and Figs. 5. It is worth noting that, having $I_{\rm L}(\phi_{\rm L}=0,\phi_{\rm R}\neq n\pi)\neq 0$ in Figs. 3(c-f) also means that a $\phi_{\rm L,0}$-Josephson junction is realized at $\Delta_{\rm M}\neq t_{\rm M}$ [74, 79, 80]. At $\phi_{\rm L}=0$, the current as a function of $\phi_{\rm R}$ can develop a $\pi$-shift for $t_{\rm M}>\Delta_{\rm M}$ [Figs. 3(e,f)], leading to a $\pi$-junction [79]. Moreover, there are also signals of a $\phi$-junction behavior [81, 82, 83, 84], where the current $I_{\rm L}$ passes through zero neither at $\phi_{\rm L}=0,2\pi$ nor $\phi_{\rm L}=\pi$ [Figs. 3(d,f)]. Yet another important effect of $\phi_{\rm L,R}\neq 0$ and $t_{\rm M}\neq\Delta_{\rm M}$ is that the maximum positive and negative values of $I_{\rm L}$ over $\phi_{\rm L}\in(0,2\pi)$ are not the same [Figs. 3(c-f)], producing a nonreciprocity between the maximum currents that can flow across the left junction. As a result, a Josephson setup with three laterally coupled minimal Kitaev chains [Eq. (2)] can realize the nonlocal Josephson effect and nonlocal nonreciprocal Josephson transport only when ECT and CAR in the middle chain are distinct.

### v2

@@ I Introduction
The research on nonlocal JE has not only attracted theoretical efforts [10, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] but also its realization has already been reported in many experiments [38, 39, 40, 41, 42, 43, 44]. Most of the studies involved conventional spin-singlet $s$-wave superconductors, leaving unconventional superconductors seldom explored in spite of their intrinsic nonlocal order parameters. One of the simplest systems that hosts controllable unconventional superconductivity with a $p$-wave order parameter can be engineered by connecting two quantum dots (QDs) via a superconductor [45, 46]. This setup represents a minimal Kitaev chain, where electron cotunneling (ECT) and crossed Andreev reflections (CARs) mediate superconducting and normal processes between QDs [45, 47]. Minimal Kitaev chains have lately triggered multiple efforts [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] and their implementation has already been shown in superconductor-semiconductor hybrids [62, 63, 64, 65]. Despite the significant attention, the nonlocal JE and its impact on minimal Kitaev chains have not yet been studied.
@@
II JJs formed by three minimal Kitaev chains
We consider a Josephson system where three minimal Kitaev chains described by
@@ II JJs formed by three minimal Kitaev chains
Here $\alpha={\rm L,M,R}$ labels the left (L), middle (M), and right (R) minimal Kitaev chains formed by two sites $j=1,2$, and $c_{\alpha_{j}}$ ($c_{\alpha_{j}}^{\dagger}$) destroys (creates) an electronic state at site $j=1,2$ in the chain $\alpha$. The first and second terms in $H_{\alpha}(\phi_{\alpha})$ describe onsite energies $\varepsilon_{\alpha_{j}}$ controlled by gates [brown bars in Fig. 1]. The third term represents hopping between sites on the same chain, a characteristic of ECT processes. The fourth term is a spinless superconducting potential, with amplitude $\Delta_{\alpha}$ and phase $\phi_{\alpha}$, describing CAR processes; note that the pair potential characterizes the $p$-wave nature of superconductivity, as in the Kitaev chain and irrespective of topology [66, 24]. We model the coupling between the three chains by the second and third terms in Eq. (2) with tunneling amplitudes $\tau_{\rm L(R)}$ across the left (right) junctions. Using a gauge symmetry, we also set the phase of the middle chain as $\phi_{\rm M}=0$, leading to the junction depending on two superconducting phases, $H(\phi_{\rm L},\phi_{\rm R})$. The phase differences $\phi_{\rm L}$ and $\phi_{\rm R}$ between chains can be externally controlled by magnetic fluxes [Fig. 1], as in Refs. [42, 67, 44]. Despite the simplicity, two-site Kitaev chains have already been experimentally realized in QDs coupled by a superconductor-semiconductor system [62, 63, 64, 65, 68], where the control of ECT and CAR was demonstrated [69, 70]. As we show below, the system exhibits a unique phase dependence, enabling a nonlocal JE with nonreciprocal origin.
III Breaking and preservation of symmetries
@@
We begin by analyzing symmetries of the Josephson system in Eq. (2). First of all, the quasi-particle energy derived by Eq. (2) forms a pair (E,E)(-E,E) because of particle-hole symmetry of the Bogoliubov-de Gennes Hamiltonian. While the Josephson system respects time-reversal symmetry (TRS), $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},-\phi_{\rm R})$, where the time-reversal anti-unitary operator $\cal T$} is given by ${\cal T}c_{\alpha_{i}}{\cal T}^{-1}=c_{\alpha_{i}}$, it does not have either $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},\phi_{\rm R})$ or $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(\phi_{\rm L},-\phi_{\rm R})$, implying local TRS is broken either across the left or right junctions, respectively. For $\tau_{\rm L(R)}\neq 0$, the left (right) junction also explicitly breaks the local spatial inversion symmetry, even when the entire system may be spatially inversion-symmetric. Thus, each junction satisfies the standard criterion for the Josephson diode effect [71, 72, 73, 74, 75].
However, our system Eq. (2) has an additional symmetry constraint for the Josephson diode effect, which originates from a special symmetry at the “sweet spot” ($\varepsilon_{\alpha}=0$, $t_{\alpha}=\Delta_{\alpha}$), where the chains support quadratically protected Majorana states [45]. In fact, for $\varepsilon_{\alpha}=0$ and $t_{\alpha}=\Delta_{\alpha}$, each minimal chain has a local charge-conjugation symmetry, ${\cal C}_{\alpha_{1}}H_{\alpha}(\phi_{\alpha}){\cal C}_{\alpha_{1}}^{-1}=H_{\alpha}(-\phi_{\alpha})$, where the non-trivial action of the unitary operator ${\cal C}_{\alpha_{1}}$ is defined by ${\cal C}_{\alpha_{1}}c_{\alpha_{1}}{\cal C}_{\alpha_{1}}^{-1}=c^{\dagger}_{\alpha_{1}}$, ${\cal C}_{\alpha_{1}}c_{\alpha_{2}}{\cal C}_{\alpha_{1}}^{-1}=-e^{i\phi_{\alpha}}c_{\alpha_{2}}$, ${\cal C}_{\alpha_{1}}c_{\beta_{i}}{\cal C}_{\alpha_{1}}^{-1}=c_{\beta_{i}}$ $(\beta\neq\alpha)$. Note that ${\cal C}_{\alpha_{1}}$ switches only an electron at site 1 on chain $\alpha$ with its antiparticle. We also have a similar local charge-conjugation symmetry ${\cal C}_{\alpha_{2}}$, switching only an electron at site 2 with its antiparticle. We can partially keep the symmetries even for the junctions in Eq. (2). For instance, by performing ${\cal C}_{{\rm L}_{2}}$, ${\cal C}_{{\rm M}_{1}}$, and an additional gauge rotation, we have ${\cal C}_{\rm LM}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm LM}=H(-\phi_{\rm L},\phi_{\rm R})$, where the unitary operator ${\cal C}_{\rm LM}$ acts as ${\cal C}_{\rm LM}c_{{\rm L}_{1}}{\cal C}^{-1}_{\rm LM}=e^{i\phi_{\rm L}}c_{{\rm L}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm L}_{2}}{\cal C}^{-1}_{\rm LM}=c^{\dagger}_{{\rm L}_{2}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{1}}{\cal C}_{\rm LM}^{-1}=-c^{\dagger}_{{\rm M}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{2}}{\cal C}_{\rm LM}^{-1}=c_{{\rm M}_{2}}$, and ${\cal C}_{\rm LM}c_{{\rm R}_{i}}{\cal C}_{\rm LM}^{-1}=c_{{\rm R}_{i}}$. We also have the relation ${\cal C}_{\rm MR}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm MR}=H(\phi_{\rm L},-\phi_{\rm R})$ with an unitary operator ${\cal C}_{\rm MR}$ constructed similarly. These relations provide the same constraint on the quasi-particle spectrum as local inversion, $E_{n}(\phi_{\rm L},\phi_{\rm R})=E_{n}(\pm\phi_{\rm L},\pm\phi_{\rm R})$, and thus, we need to break these local charge-conjugation symmetries, in addition to local TRS, to obtain the Josephson diode effect. For this purpose, we consider $\Delta_{\rm M}\neq t_{\rm M}$ below, which breaks both ${\cal C}_{\rm LM}$ and ${\cal C}_{\rm MR}$.
@@ IV Andreev bound states and Andreev molecules
Having identified the role of the superconducting phase differences $\phi_{\rm L,R}$ on the system symmetries, here we explore how they determine the behavior of the Andreev energy spectrum. This is shown in Fig. 2(a-d) as a function of $\phi_{\rm L}$ when ECT and CAR in the middle chain are the same ($t_{\rm M}=\Delta_{\rm M}$) for distinct values of $\phi_{\rm R}$ at $\Delta_{\rm L,R}=t_{\rm L,R}$, while in Fig. 2(e-h) for ECT weaker than CAR. Since $\Delta_{\rm L,R}=t_{\rm L,R}$, the left (right) chains host zero-energy dispersionless Majorana flat bands located in their first (second) QD, which remain robust against $\phi_{\rm L,R}$ [Fig. 2]. Depending on the ratio between CAR and ECT in the middle chain, the Andreev spectrum exhibits a strong dependence on $\phi_{\rm L,R}$. For $t_{\rm M}=\Delta_{\rm M}$ at $\phi_{\rm R}=0$ [Fig. 2(a)], a pair of dispersionless Andreev bound states (ABSs) appear at energies close to $\pm\Delta$ and intersect a pair of dispersing ABSs near $\phi_{\rm L}=0,2\pi$. See green vertical lines. These dispersionless ABSs originate from the right junction between the middle and right chains, and, as $\phi_{\rm R}$ takes finite values, they hybridize with the dispersing ABSs formed in the left junction between the left and middle chains [Figs. 2(a-d)]. The dispersing hybridized ABSs, which can be seen as Andreev molecules studied before in conventional superconductors [10, 76, 77, 26, 78, 43, 79, 67, 32], reach zero energy at $\phi_{\rm L}=\pi$ (gray vertical line) and become flat bands with $\phi_{\rm L}$ at $\phi_{\rm R}=\pi$ [Fig. 2(c)], always maintaining a symmetric profile about $\phi_{\rm L}=\pi$.
When $t_{\rm M}\neq\Delta_{\rm M}$, a nonzero $\phi_{\rm R}$ modifies the phase dependence of the spectrum and, notably, it also makes the spectrum asymmetric about $\phi_{\rm L}=\pi$ (vertical dashed red line), provided $\phi_{\rm R}\neq 2\pi n$ with $n\in\mathbb{Z}$; see Fig. 2(e-h) and Fig. 5 in Appendix A. This asymmetry moves the zero-energy crossing of the dispersing hybridized ABSs to $\phi_{\rm L}\neq\pi$, which occurs below (above) $\phi_{\rm L}=\pi$ when $\phi_{\rm R}<\pi$ ($\phi_{\rm R}>\pi$) for $t_{\rm M}<\Delta_{\rm M}$ [Fig. 2(e-h)]. The asymmetric spectrum turns out to be gapless over $\phi_{\rm L,R}\in(0,2\pi)$ and remains for $t_{\rm M}>\Delta_{\rm M}$ but with the zero-energy crossing at $\phi_{\rm L}<\pi$ ($\phi_{\rm L}>\pi$) when $\phi_{\rm R}>\pi$ ($\phi_{\rm R}<\pi$) in the latter case [Fig. 5]. A consequence of the asymmetric spectrum under $\phi_{\rm L,R}\neq 0$ and $\Delta_{\rm M}\neq t_{\rm M}$ is that the free energy develops minima away from $\phi_{\rm L}=2\pi n$, unlike conventional Josephson junctions [80, 81]; (See Appendix A). Thus, the imbalance between CAR and ECT in the middle Kitaev chain induces an asymmetric Andreev spectrum at nonzero superconducting phases across the left/right junctions, in line with breaking local TRS and local charge-conjugation symmetry. From an experimental point of view, we note that the ratio between $\Delta_{\rm M}$ and $t_{\rm M}$ can be controlled e. g., by gating the middle chain as done in Ref. [62] for tuning CAR and ECT processes.
@@ V Local and nonlocal Josephson effect
A direct consequence of $\phi_{\rm L,R}\neq 0$ and the imbalance between CAR and ECT in the middle chain is the phase-biased Josephson effect, which we explore here. For this purpose, we numerically obtain the phase-biased Josephson current flowing across the left junction as [82, 83, 24] $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm L}$  [^1] Similarly, we can obtain the supercurrent flowing across the right junction as $I_{\rm R}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm R}$. We have verified that $I_{\rm L,R}(\phi_{\rm L},\phi_{\rm R})$ can also be obtained within a Green’s function approach [109, 83], but obtaining them from the spectrum offers a direct way to identify the contributing states [110, 111, 112, 113, 114, 115, 116].. To visualize the supercurrent, in Figs. 3(a,c,e) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ and $\phi_{\rm R}$ for distinct cases when CAR in the middle chain is equal (a), weaker (c), and stronger (e) than ECT in the same chain. In Figs. 3(b,d,f) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ for distinct $\phi_{\rm R}$, while Figs. 3(d,f) also show $I_{\rm L}$ as a function of $\phi_{\rm R}$ for distinct $\phi_{\rm L}$. The first observation is that $I_{\rm L}$ in Figs. 3 strongly depends on $\phi_{\rm L,R}$ and the ratio between ECT ($t_{\rm M}$) and CAR ($\Delta_{\rm M}$) in the middle region, a behavior that originates from the ABSs in Figs. 2. When $\Delta_{\rm M}=t_{\rm M}$, the supercurrent $I_{\rm L}$ is $2\pi$-periodic and symmetric under $\phi_{\rm R}\neq 0$ [Figs. 3(a,b)], having $I_{\rm L}(\phi_{\rm L}=n\pi)=0$ for any $\phi_{\rm R}$ and equal maxima/minima that reduce for $\phi_{\rm R}\neq 0,2\pi$; the free energy in Fig. 6(a) of Appendix B has a minimum at $\phi_{\rm L}$ as for 0-junctions [80, 85]. Also, $I_{\rm L}$ exhibits a sawtooth profile at $\phi_{\rm L}=\pi$ for $\phi_{\rm R}\neq\pi$ due to the zero-energy crossing of the dispersing ABSs [Figs. 2(a-d)], while it smoothens out at $\phi_{\rm R}=\pi$ because the dispersing ABSs become flat [Figs. 2(c)].
@@
At $\Delta_{\rm M}\neq t_{\rm M}$, the supercurrent across the left junction $I_{\rm L}$ develops an intriguing anomalous behavior that varies with $\phi_{\rm L,R}\neq 0$. See Figs. 3(c-f). The most notorious feature is that $\phi_{\rm R}\neq n\pi$ gives rise to $I_{\rm L}(\phi_{\rm L}=n\pi)\neq 0$ with $n\in\mathbb{Z}$ and it is not symmetric about $\phi_{\rm L}=\pi$ anymore, which are consequences of the asymmetric Andreev spectrum shown in Figs. 2(e-h). This implies that, even at $\phi_{\rm L}=0$, a supercurrent across the left junction $I_{\rm L}$ emerges and is controlled by the superconducting phase difference across the right junction $\phi_{\rm R}$, signalling the so-called nonlocal JE which was initially predicted in systems with conventional superconductors [10]. This nonlocal JE requires multiple terminal and does not appear in a two-terminal Josephson junction. The sawtooth profile of $I_{\rm L}$ also occurs away from $\phi_{\rm L}=\pi$ [Figs. 3(d,f)] due to the shifted zero-energy crossing in the ABSs when $t_{\rm M}\neq\Delta_{\rm M}$. See Figs. 2(e-h) and Figs. 5. It is worth noting that, having $I_{\rm L}(\phi_{\rm L}=0,\phi_{\rm R}\neq n\pi)\neq 0$ in Figs. 3(c-f) also means that a $\phi_{\rm L,0}$-Josephson junction is realized at $\Delta_{\rm M}\neq t_{\rm M}$ [80, 85, 86]. At $\phi_{\rm L}=0$, the current as a function of $\phi_{\rm R}$ can develop a $\pi$-shift for $t_{\rm M}>\Delta_{\rm M}$ [Figs. 3(e,f)], leading to a $\pi$-junction [85, 87]. Moreover, there are also signals of a $\phi$-junction behavior [88, 89, 90, 91], where the current $I_{\rm L}$ passes through zero neither at $\phi_{\rm L}=0,2\pi$ nor $\phi_{\rm L}=\pi$ [Figs. 3(d,f)]. Yet another important effect of $\phi_{\rm L,R}\neq 0$ and $t_{\rm M}\neq\Delta_{\rm M}$ is that the maximum positive and negative values of $I_{\rm L}$ over $\phi_{\rm L}\in(0,2\pi)$ are not the same [Figs. 3(c-f)], producing a nonreciprocity between the maximum currents that can flow across the left junction. As a result, a Josephson setup with three laterally coupled minimal Kitaev chains [Eq. (2)] can realize the nonlocal Josephson effect and nonlocal nonreciprocal Josephson transport only when ECT and CAR in the middle chain are distinct.
@@
VI Critical currents and nonlocal Josephson diode effect